OpenStudy (anonymous):

Ive been sick for a couple of weeks please help! Harriet needs to ship a small vase. The box she will use has a volume of 216 cubic inches. If the side lengths are all the same, what is the length of each side of the box? Hint: V = s3

OpenStudy (chihiroasleaf):

V = s^3 216 = s^3 so, what is s? :)

OpenStudy (anonymous):

i missed the lesson, so idk

OpenStudy (anonymous):



OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

216=6x^2 x^2=216/6 x=\[\sqrt{216/6}\]

OpenStudy (anonymous):

ok